If a+b+c=1 and ab+bc+ca=2 and abc=3 then (1/a+bc)+(1/b+ac)+(1/c+ab)=?

Grade:9

2 Answers

Vijay Mukati
7 years ago
Hint:

Take the LCM of the given equation and use the given results to simplify and get its value.

Thanks, Vj
Kumud Garain
2 years ago
a+b+c=1 ab+bc+ca=2 abc=3 (1)/(a+b)+(1)/(b+ca)+(1)/(b+ca)=? Find the above equation. Sir please.        Do solve this problem